OpenStudy (unicornpoopcookies):

If the sum of interior angle measures of a polygon is 900 degrees how many sides does the polygon have?

OpenStudy (anonymous):

number of sides and interior angle of polygons is related by the following formula \[interior \ \angle=(number\ of \ sides-2)*180\] hope that helps!

OpenStudy (unicornpoopcookies):

900 = 180 (s-2) 5 = s-2 7 = s Is this right?

OpenStudy (anonymous):

yep!